A.) |2x + 2| = 14
x=
B.) |x| = 6
x=
C.) -5|x + 1| = -10, the isolated absolute value equation would be |x + 1| = -15.
I think its true...can you help me out?
D.) 2|x + 4| = 8, the isolated absolute value equation would be |x + 4| = 4.
I think its false...What do you think?
E.)|x – 2| + 10 = 12
X=
F.)|2x + 6| - 4 = 20
x =
G.)6 – 3|2x + 6| = 0
X=

H.) Solve for x. 10 - |x + 2| = 12


I.) Solve for x. -5|x + 1| = -10

Answer by stanbon(48546)

A.) |2x + 2| = 14
2x+2 = 14 or 2x+2 = -14
x+1 = 7 or x+1 = -7
x = 6 or x = -8
-------------------------
B.) |x| = 6
x=6 or x = -6
-------------------------

C.) -5|x + 1| = -10
|x+1| = 2
x+1 = 2 or x+1 = -2
x = 1 or x = -3
--------------------------
D.) 2|x + 4| = 8
|x+4| = 4
x+4 = 4 or x+4=-4
x = 0 or x = -8
-----------------------------
E.)|x – 2| + 10 = 12
|x-2| = 2
x-2= 2 or x-2=-2
x = 4 or x = 0
--------------------------

F.)|2x + 6| - 4 = 20
|2x+6| = 24
2x+6 = 24 or 2x+6 = -24
2x = 18 or 2x = -30
x = 9 or x = -15
--------------------------

G.)6 – 3|2x + 6| = 0
3|2x+6| = 6
2x + 6 = 2 or 2x+6 = -2
2x = -4 or 2x = -8
x = -2 or x = -4
-------------------------
H.) 10 - |x + 2| = 12
|x+2} = -2
No solution as absolute cannot equal a negative value.
==========================================================


I.) -5|x + 1| = -10
|x+1| = 2
x+1 = 2 or x+1 = -2
x = 1 or x = -3
======================
Cheers,
Stan H.